OpenStudy (anonymous):

Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770?

OpenStudy (anonymous):

suppose adults are X and students are Y total students are 150 this means X+Y=150 Y=150-X-------(eq 1) now, 45.00X+28.50Y=4770 45.00X+28.50(150-X)=4770 45.00X+4275-28.50X=4770 16.50X=4770-4275 16.50X=495 X=495/16.50 X=30 so we have 30 adults now go back to eq 1 Y=150-X Y=150-30 Y=120 so we have 120 students!! Hope it helps!